# MotionAgent: Fine-grained Controllable Video Generation via Motion Field Agent

### v1

@@ 4.2.1 General I2V Generation
Compared with the base model SVD [4], our method shows better results in consistency, smoothness, and aesthetic quality metrics. These improvements are attributed to the motion field agent, which can reason and generate suitable object trajectories and camera motion even when there is no motion information included in the text prompts. Additionally, the tuning of the optical flow adapter also improves video quality by constraining the generated video with optical flow and eliminating unreasonable motion. The dynamic degree metric of our method decreases due to the precise control we implement over the generated video. The objects mentioned in the text move accurately, while those not mentioned remain as static as possible, which leads to a lower dynamic degree score. When our model is given a detailed text description with more motion information, it can generate videos with a higher dynamic degree.
@@ 4.2.2 Controllable I2V Generation
We reuse images from VBench [20] and add more motion information to the original text prompts. To evaluate the control capabilities of object movement, we design new text prompts for 83 images from VBench, providing detailed object movement descriptions. For camera motion, we modify the simple camera motion prompts of 109 images to more complex ones, such as changing “zoom in” to “first zoom in then zoom out” or “zoom in incrementally.”
@@
MotionAgent (Ours) 45.69 77.76 65.10
Table 2: Evaluation results of controllable I2V generation on our benchmark (all values are in percentage). The best result is indicated in bold, the second-best result is indicated with underlines, and the third-best result is indicated with double underlines.
As shown in Table 2, our method achieves the best results compared to other methods in controllable video generation tasks. For the Object Movement Q&A metric, our method demonstrates stronger control over object movement, with higher accuracy scored by GPT-4o [34]. It is worth noting that current multimodal LLMs have limited reasoning capabilities, which may lead to some inaccurate judgments and responses. However, the numerical results still indicate that our method achieves the most accurate control over object movement. For the Complex Camera Motion metric, our method significantly outperforms the other methods. Compared to Motion-I2V [39], the optical flow generated by the motion field agent represents the motion information more precisely and results in a higher total score. Moreover, our method maintains consistent performance regardless of whether the camera motion is simple or complex, as illustrated by the camera motion metrics in Tables 1 and 2. Some qualitative comparison results are shown in Figure 4, where our method showcases more precise control over generated videos.
Our method directly converts motion information in the text prompts into the intermediate motion representations through the motion field agent, and then uses optical flow to achieve controllable video generation. Without needing high-quality video-image-text pair training data, our approach achieves fine-grained controllable I2V generation.
@@ 4.2.3 User Study
We collect 30 images from VBench [20] and expand the prompts by adding detailed descriptions of object movement and camera motion. We then generate videos using the official codes of DynamiCrafter [54], CogVideoX [60], and Pyramid Flow [21]. The user study is expected to be completed in 10-20 minutes. For each case, the user study interface shows the videos generated by four methods, and participants are instructed to evaluate the videos from two dimensions: (i) “Please sort the videos by visual quality from best to worst.”; (ii) “Please sort the videos from high to low based on the semantic alignment of the motion information in the text prompt and the video.”. Finally, we receive 50 valid responses from the participants.
@@ 4.3 Ablation Study
Detection Object Camera Flow Object Movement Complex Dynamic
Tool Movement Motion Composition Q&A Camera Motion Degree
34.33 75.11 20.53
38.20 0.30 29.95
10.51 75.95 8.42
30.07 64.92 27.89
45.69 77.76 32.11
Table 3: Ablation study of object identification module and optical flow composition module on our benchmark (all values are in percentage).
@@ 4.3.1 Object Identification
We compare the two approaches for identifying the objects described in the text. As shown in Table 3, the results using Grounded-SAM (full model) are significantly better than those using the multi-round dialogue (without a detection tool). The multi-round dialogue leads to some incorrect judgments, resulting in a lower score on the Object Movement Q&A metric. Additionally, our method decomposes object movement and camera motion, so the degradation caused by incorrect object identification does not significantly affect the camera motion in the generated video. Moreover, the misjudgments in object identification lead to the degradation of the dynamic degree metric.
4.3.2 Optical Composition
@@
Then, we conduct an ablation study on the analytical optical flow composition module and use the original generative motion field adapter proposed by MOFA-Video [32] as a baseline. The baseline can only control object movement or camera motion independently, so we define two models: “without camera motion” and “without object movement,” which represent models that can only control either object movement or camera motion independently.
In our proposed module, we design an approach to compose optical flow using an analytical method. To verify the effectiveness of this approach, we design a mechanism that directly adds the optical flow of object movement and the optical flow caused by camera motion, which we call “without flow composition.”
@@
As shown in Table 3, the baseline models can only control object movement or camera motion independently. Compared to “without flow composition,” our proposed analytical optical flow composition module achieves higher scores on both the Object Movement Q&A and Complex Camera Motion metrics. Directly adding optical flows results in many unreasonable areas in the added flow maps, and the object optical flow may be overshadowed by the camera optical flow. For the dynamic degree metric, our full module also achieves the highest scores. Moreover, the dynamic degree results are higher than those acquired from the original prompts in VBench [20]. This showcases that our method generates videos with a higher dynamic degree when there is more detailed motion information in the given text, and follows the text instructions accurately.
@@ 5 Conclusion
This work aims to address the challenge of precisely controlling I2V generation using only text input. We first propose a motion field agent that converts the motion information in the text into object trajectories and camera extrinsics. Then, we design an analytical optical flow composition module to integrate the explicit motion representations into unified optical flow maps. Next, we fine-tune an optical flow adapter to achieve controllable I2V generation. In the experiments, we construct a new video motion benchmark, and the evaluation results demonstrate that our method significantly outperforms other text-guided I2V generation models.
Limitation. Due to the limited reasoning capabilities of LLMs, our method may fail to understand particularly complex motions. In these cases, the object trajectories and camera extrinsics generated by the motion field agent may be unreasonable. In future work, we can address these problems by constructing text-motion pairs and fine-tuning the LLMs to enhance their ability to understand particularly complex motions.
Supplementary Material
6 Object Trajectory Plotting Module
6.1 Details of Trajectory Plotting
Inspired by AppAgent [65], we replace direct trajectory generation with grid selection based on grid numbers. Specifically, we divide the given image into $N\times M$ grids, breaking it down into small square areas. Each area is labeled with an integer in the top-left corner and subdivided into nine subareas. Based on the previous step, we identify the starting point of the trajectory using the detection result and plot this starting point on the image, represented by a circle. Then, we provide the image overlaid with the grids and starting point as input to the agent.
Figure 6: Details of trajectory plotting: The grid divides the image into small square areas. Each area is labeled with an integer in the top-left corner and is further subdivided into nine subareas.
We define the following functions: Set_*_Points (start: int, string; mid_*: int, string; end: int, string) for the agent. Here, * is an integer ranging from $1-4$, used to achieve varying lengths and curvature in trajectory plotting. The parameters start, mid_*, and end include an integer label assigned to the grid area and a string representing the exact location within the grid area. The string can take one of the following nine values: center, top-left, top, top-right, left, right, bottom-left, bottom, and bottom-right.
A simple use case is Set_2_Points (start: 143, top-right; end: 33, bottom-right), which sets the starting point of the trajectory at the top-right of grid area 143 and the endpoint at the bottom-right of grid area 33. As illustrated in Figure 6, this function represents a simple linear trajectory.
Once we obtain the complete object trajectory, we use interpolation to determine the position in the trajectory for each frame. Then, these interpolated positions are input into the subsequent model to calculate dense optical flow.
Offset Grid Object Movement Dynamic
Generation Selection Q&A Degree
29.28 7.63
45.69 32.11
Table 4: Ablation study of trajectory plotting module (all values are in percentage).
6.2 More Ablation Study
Here, we conduct an ablation study on different approaches for trajectory plotting. We use direct offset generation as the baseline module instead of grid selection. Specifically, we provide the agent with the starting point location based on the detection results and ask the agent to directly generate offsets from the starting point to define the object trajectory.
As shown in Table 4, the grid selection approach shows better evaluation results in both metrics. The grid selection approach gives the agent an overall understanding of the image layout, and it is easier for the agent to use than direct offset generation. As for the dynamic degree metrics, the offset generation approach cannot output a suitable trajectory length, which may lead to a lower dynamic degree.
Figure 7: Intermediate representations generation by motion field agent, including object trajectory, camera extrinsics, and unified optical flow maps.
7 Training Data Preparation
To eliminate the domain gap between the unified and real optical flow maps, we propose fine-tuning the optical flow adapter module, which maintains the generation capabilities of the base I2V diffusion model. For each video used for training, we first utilize a binary segmentation model, BiRefNet [70], to decompose the foreground and background. We then remove the dynamic foreground based on the binary segmentation mask. Next, we adopt an SLAM method, DROID-SLAM [46], to compute the camera extrinsics $\hat{E}$ for each frame based on the masked video. Additionally, for the original video, we use an optical flow model, Unimatch [57], to estimate the real optical flow $\hat{F}$.
Next, we explain how to estimate the optical flow caused by object movement based on the camera extrinsics $\hat{E}$ and real optical flow $\hat{F}$. We define $I^{0}$ as the pixel position in the first frame. According to the predicted real optical flow $\hat{F}$, we compute the corresponding pixel position in the following frames, which can be formulated as:
$\hat{I}=I^{0}+\hat{F}.$ (5)
We reproject the pixel position in the following frames back to the first frame based on the predicted camera extrinsics $\hat{E}$. This can be computed by:
$\hat{I}^{0}=\Pi(\hat{E}^{-1}\Pi^{-1}(\hat{I})),$ (6)
where $\Pi$ and $\Pi^{-1}$ are the projection and unprojection operations, respectively. $\hat{I}^{0}$ indicates the pixel position in the first frame that contains only object movement. Finally, we compute the optical flow caused by object movement:
$\hat{F}_{obj}=\hat{I}^{0}-I^{0}.$ (7)
We perform sampling [64] on these optical flow maps of object movement to obtain sparse object trajectories. Subsequently, we reuse the proposed analytical composition method to calculate the unified optical flow maps $\bar{F}$, which are utilized as input to fine-tune the optical flow adapter. Additionally, we calculate the error between the unified optical flow maps $\bar{F}$ and the real optical flow $\hat{F}$. If the error exceeds a threshold, we replace the unified optical flow maps $\bar{F}$ with the real optical flow $\hat{F}$ for training.
8 Qualitative Results
8.1 Visualization of Intermediate Representations
As illustrated in Figure 7, we show the intermediate representations generation by the motion field agent.
8.2 More Comparison Results
In Figure 8, we show more results compared to DynamiCrafter [54], CogVideoX [60] and Pyramid Flow [21].
9 User Study Interface
Figure 9 shows the designed user study interface. For each question, we randomly shuffle four videos generated by our method and the other three methods. We then ask participants to rank the videos from highest to lowest twice based on specific requirements. After the user study, we calculate the mean ranking for each method across different evaluation dimensions.
Figure 8: More comparison results of controllable I2V generation on our benchmark. The motion described in the text is in bold.
Figure 9: User study interface. Each participant is required to evaluate 30 groups of videos and respond to two corresponding sub-questions for each group. Due to the page limit, only one group of videos and two sub-questions are shown here.

### v2

@@ 4.2.1 General I2V Generation
Compared with the base model SVD [5], our method shows better results in consistency, smoothness, and aesthetic quality metrics. These improvements are attributed to the motion field agent, which can reason and generate suitable object trajectories and camera motion even when there is no motion information included in the text prompts. Additionally, the tuning of the optical flow adapter also improves video quality by constraining the generated video with optical flow and eliminating unreasonable motion. The dynamic degree metric of our method decreases due to the precise control we implement over the generated video. The objects mentioned in the text move accurately, while those not mentioned remain as static as possible, which leads to a lower dynamic degree score. When our model is given a detailed text description with more motion information, it can generate videos with a higher dynamic degree. The qualitative results of dynamic degree metric can be found in the supplementary materials.
@@
Figure 5: Fine-grained controllable video results generated by our method. Lines 1-2 are multiple object movements control; Line 3 is camera motion control.
@@ 4.2.2 Controllable I2V Generation
We reuse images from VBench [26] and add more motion information to the original text prompts. To evaluate the control capabilities of object movement, we design 432 new text prompts for 83 images from VBench, providing detailed object movement descriptions. For camera motion, we modify the simple camera motion prompts from 109 images to 662 more complex prompts, such as changing “zoom in” to “first zoom in then zoom out” or “zoom in incrementally.”
@@
MotionAgent 45.69 77.76 65.10
MotionAgent (Rethinking) 49.58 89.04 73.45
Table 2: Evaluation results of controllable I2V generation on our benchmark (all values are in percentage). The best result (before rethinking) is indicated in bold, the second-best result is indicated with underlines, and the third-best result is indicated with double underlines.
As shown in Table 2, our method achieves the best results compared to other methods in controllable video generation tasks. For the Object Movement Q&A metric, our method demonstrates stronger control over object movement, with higher accuracy scored by GPT-4o [41]. It is worth noting that current multimodal LLMs have limited reasoning capabilities, which may lead to some inaccurate judgments and responses. However, the numerical results still indicate that our method achieves the most accurate control over object movement. For the Complex Camera Motion metric, our method significantly outperforms the other methods. Compared to Motion-I2V [46], the optical flow generated by our motion field agent represents the motion information more precisely and results in a higher total score. Moreover, our method maintains consistent performance regardless of whether the camera motion is simple or complex, as illustrated by the camera motion metrics in Tables 1 and 2. Some qualitative comparison results are shown in Figure 4, where our method showcases more precise control over generated videos.
Our method directly converts motion information in the text prompts into the intermediate motion representations through the motion field agent, and then uses optical flow to achieve controllable video generation. Without the need for high-quality video-image-text pair training data, our approach achieves fine-grained controllable I2V generation. Figure 5 illustrates some Fine-grained controllable video results, and our method achieves precise control over multiple object movements and sophisticated camera motion.
@@ 4.2.3 User Study
We collect 30 images from VBench [26] and expand the prompts by adding detailed descriptions of object movement and camera motion. We then generate videos using the official codes of DynamiCrafter [63], CogVideoX [68], and Pyramid Flow [27]. The user study is expected to be completed in 10-20 minutes. For each case, the user study interface shows the videos generated by four methods, and participants are instructed to evaluate the videos from two dimensions: (i) “Please sort the videos by visual quality from best to worst”; (ii) “Please sort the videos from high to low based on the semantic alignment of the motion information in the text prompt and the video”. Each case is rated by multiple participants. Finally, we receive 50 valid responses and the final results are calculated based on the 1500 ratings.
@@ 4.3.1 Object Identification
We compare the two approaches for identifying the objects described in the text. As shown in Table 3, the results using Grounded-SAM (full model) are significantly better than those using the multi-round dialogue (without a detection tool). The multiround dialogue leads to some incorrect judgments, resulting in a lower score on the Object Movement Q&A metric. It is worth mentioning that our method decomposes object movement and camera motion, so the degradation caused by incorrect object identification does not significantly affect camera motion in the generated video. Moreover, misjudgments in object identification lead to degradation of the dynamic degree metric.
Detection Object Camera Flow Adapter Object Movement Complex Dynamic
Tool Movement Motion Composition Fine-tuning Q&A Camera Motion Degree
34.33 75.11 20.53
10.51 75.95 8.42
38.20 0.30 29.95
30.07 64.92 27.89
40.56 48.27 28.47
45.69 77.76 32.11
Table 3: Ablation study of object identification module, optical flow composition module and adapter tuning on our benchmark (all values are in percentage).
Figure 7: Comparison of before and after rethinking process.
4.3.2 Optical Composition and Adapter Fine-tuning
@@
We first conduct an ablation study on the analytical optical flow composition module and use the original generative motion field adapter proposed by MOFA-Video [39] as a baseline. The baseline can only control object movement or camera motion independently, so we define two models: “without camera motion” and “without object movement”, representing models that can only control object movement or camera motion independently. In our proposed module, we design an analytical approach to compose optical flow. To verify the effectiveness of this approach, we design a mechanism that directly adds the optical flow of object movement and the optical flow caused by camera motion, which we call “without flow composition”.
@@
As shown in Table 3, the baseline models can only control object movement or camera motion independently. Compared to “without flow composition”, our proposed analytical optical flow composition module achieves higher scores on both the Object Movement Q&A and Complex Camera Motion metrics. Directly adding optical flows results in many unreasonable areas in the added flow maps, and the object optical flow may be overshadowed by the camera optical flow.
Then, we conduct an ablation study on the effectiveness of optical flow adapter tuning. In Table 3, the primitive optical flow adapter is only well suited to optical flow caused by the object movement. When we input unified optical flow maps containing motion information from both object and camera movement, the primitive adapter, without tuning, produces poor results on complex camera motion.
For the dynamic degree metric, our full module achieves the highest scores. Moreover, the results of the dynamic degrees are higher than those acquired from the original prompts in VBench [26]. This showcases that our method generates videos with a higher dynamic degree when there is more detailed motion information in the given text, and follows the text instructions accurately.
Method Object Movement Complex Total
Q&A Camera Motion Scores
MotionAgent(w Qwen2 [3] & SVD [5]) 45.07 72.62 61.73
MotionAgent(w Llama3 [12] & SVD [5]) 45.63 75.66 63.80
MotionAgent(w GPT-4o [1] & Motion-I2V [46]) 44.57 71.73 61.00
MotionAgent(w GPT-4o [1] & SVD [5]) 45.69 77.76 65.10
Table 4: Ablation study results on different LLMs and base I2V generation model on our benchmark (all values are in percentage).
4.3.3 Different LLMs and Base I2V Generation model
As illustrated in Table 4, we further replace GPT-4o [1] with other open-sourced LLMs (QWen2 [3] and LLama3 [12]) to verify the robustness of our method. These alternative LLMs achieve almost the same performance, demonstrating that our method is robust with both closed-sourced and open-sourced LLMs. Besides, we also utilize Motion-I2V [46] as an additional baseline I2V model to demonstrate the generalization of our method. Specifically, we replace Flow U-Net of Motion-I2V with our method and input the unified optical flow maps into I2V U-Net.
4.3.4 Rethinking
In Table 2, we report the evaluation results before and after the rethinking step on our benchmark. After the rethinking progress, our method improves the metrics of complex camera motion obviously. By analyzing the generated video, our method identifies the camera motion that is not precise enough and generates a more suitable camera path. Moreover, this step also refines the plotted object trajectories, which the metric of object movement can demonstrate. Figure 7 shows an example of the rethinking steps, the agent amplifies the translation in the z-axis of the camera motion according to the generated video.
@@ 5 Conclusion
This work aims to address the challenge of precisely controlling I2V generation using only text input. We first propose a motion field agent that converts the motion information in the text into object trajectories and camera extrinsics. Then, we design an analytical optical flow composition module to integrate the explicit motion representations into unified optical flow maps. Next, we fine-tune an optical flow adapter to achieve controllable I2V generation. Finally, we introduce a rethinking mechanism that enables the agent to correct previous actions according to the generated results. In the experiments, we construct a new video motion benchmark, and the evaluation results demonstrate that our method significantly outperforms other text-guided I2V generation models.
6 Acknowledgments
This research is supported by the MoE AcRF Tier 2 grant (MOE-T2EP20223-0001).